OpenStudy (anonymous):

A ball is thrown downward at 5ms from roof 10m high its velocity when it reaches the ground is?

OpenStudy (anonymous):

5x10x9.8(speed of gravity)

OpenStudy (anonymous):

the answer is 15m/s, but i dont know how to get it.

OpenStudy (anonymous):

The movement equation is \[V=V _{0}+g t\], where \[V _0\] is the initial velocity, in your case 5 m/s, g the gravity acceleration \[9,8 m/s ^{2}\] (you can use the aprox. 10 m/s2, I do not know if it is important for your calculations). Know, you need to know how much time does the ball need to reach the ground. Apply this equation, \[x=x_0-V_0t\] With this you know the time. Introduce it in the first equation to know the final velocity in the time you calculated. Well, a final note, there no succh thinh as the speed of gravity. Acceleration is sommething different from velocity, its the time rate of it It help? feel free to ask if didn't understand sommething. Bye. António

OpenStudy (anonymous):

fqantonio I really liked your explanation. I'd like to add, that there is another equation that would fit this problem nicely though. Vyf^2 = Vyi^2 + 2 * ay * (yf-yi) where Vyf is final velocity in respect to y Vyi is initial velocity in respect to y ay is acceleration in respect to y yf is final y position yi is initial y position. to find Vyf all you do is get the square root of the right side. No need for time in this equation.

OpenStudy (anonymous):

use v=u + a*t to calculate the velocity. In ordre to find the value of t: s=u*t + (a*t^2)/2 here s=10 a=9.8 u=5 after calculation u get t=1,-4. since -4 is not valid.consider 1. now substitute t=1 in the first equation. u get v=14.8m/s as said by......